# A rank decomposition for the topological classification of neural representations

### v1

@@ Abstract
Neural networks can be thought of as applying a transformation to an input dataset. The way in which they change the topology of such a dataset often holds practical significance for many tasks, particularly those demanding non-homeomorphic mappings for optimal solutions, such as classification problems. In this work, we leverage the fact that neural networks are equivalent to continuous piecewise-affine maps, whose rank can be used to pinpoint regions in the input space that undergo non-homeomorphic transformations, leading to alterations in the topological structure of the input dataset. Our approach enables us to make use of the relative homology sequence, with which one can study the homology groups of the quotient of a manifold $\mathcal{M}$ and a subset $A$, assuming some minimal properties on these spaces.
As a proof of principle, we empirically investigate the presence of low-rank (topology-changing) affine maps as a function of network width and mean weight. We show that in randomly initialized narrow networks, there will be regions in which the (co)homology groups of a data manifold can change. As the width increases, the homology groups of the input manifold become more likely to be preserved. We end this part of our work by constructing highly non-random wide networks that do not have this property and relating this non-random regime to Dale’s principle, which is a defining characteristic of biological neural networks. Finally, we study simple feedforward networks trained on MNIST, as well as on toy classification and regression tasks, and show that networks manipulate the topology of data differently depending on the continuity of the task they are trained on.
@@ 1 Introduction
Following this, we introduce the idea of a representation classifying space of an input manifold, which makes use of the relative homology groups of the input manifold with respect to subspaces that fall in the low-rank regions of the polytope decomposition. We analytically compute some examples with very simple manifolds. This is followed up by empirically studying randomly initialized narrow and wide neural networks, for which we provide conditions that determine whether such networks will implement either homeomorphic or non-homeomorphic maps. We also explore the impact on input topology in networks following variations of Dale’s principle. Finally, we explore which regime networks trained to solve regression and classification tasks converge to. We hope that this work provides an interesting perspective and opens up the path towards the application of more advanced topological methods in the study of neural networks.
@@ 2.1 Neural networks and decompositions
In this equation $J$ is a binary vector denoting the neurons in layer $K$ which are active and $S(x)$ is a vector function. The quantity $||J||_{0}$ will be referred to as the codeword size. These codewords form a decomposition of the input space at each layer of the network $\mathcal{M}=\bigsqcup\limits_{J}C_{J}^{K}$ for any $K$. We can also imagine a more global decomposition which is not layer dependent by defining $J$ over all the neurons in the network instead of layerwise. Following the vocabulary in [8], we call that the global polytope decomposition $C_{J}$. Another feature of this decomposition is that within each codeword set, we can express the map $\Phi^{K}$ as an affine map $Q_{J_{K}}(W_{K}\Phi^{K-1}(x)+b_{K})$, where $Q_{J_{K}}$ is a matrix which is 0 everywhere except on the diagonal $\text{diag}(Q_{J_{K}})=J_{K}$. In a way it essentially substitues the ReLU function at layer $K$ for a piecewise affine map. As one can see this definition is recursive and by going backwards through the first layer, it can be written as,
@@ 2.2 Non-homeomorphic regions of neural manifolds
Since full-rank affine maps between vector spaces are continuous and invertible bijections, they induce a homeomorphism between manifolds embedded in these spaces and are therefore not able to change the topology of a manifold. The only case in which the topology can change is when such an affine map projects subsets of the manifold to a subspace of a lower dimension than its embedding dimension. This is due to the fact that the projected subsets can no longer be bijectively embedded in this new lower-dimensional space. Such a degenerate map is non-injective and therefore non-homeomorphic. In order to classify the types of manifolds that one can get through such mappings, we focus on studying a standard topological invariant, namely (co)homology groups [9]. Assuming one is working with nice (compact and Hausdorff spaces as shown in theorem 2.14 in [29]) input spaces, when the map $\Phi$ between manifolds fails to be injective, studying the homology groups $H(\Phi(\mathcal{M}))$ is equivalent to studying $H(\mathcal{M}/\sim_{\Phi})$, where $\sim_{\Phi}$ contains the equivalence classes induced by the equivalence relation $\Phi(x)=\Phi(y)$ and $\mathcal{M}/\sim_{\Phi}$ is equipped with the quotient topology.
Since the mapping $\Phi$ is a homeomorphism on all regions $\Omega_{n}$ of rank equal to $m$, we only need to consider what happens to regions of lower rank. Let $\Omega_{0}^{K}$ be a rank 0 subset from the input manifold which is projected to a single point. If this is a contractible space, projecting it to a point will not change any of the (co)homology groups. Otherwise, quotiening this space out can have an impact. Both of these cases can be studied through the relative homology sequence [9], which is defined for a pair of spaces $(X,A)$ with $A\subset X$. If one has a map which is homeomorphic on all points of $\mathcal{M}$ except on $\Omega_{0}^{K}$ which is all projected to a point, then the relative homology sequence of the pair $(\mathcal{M},\Omega_{n}^{K})$ can be studied. Since $\Phi$ is not necessarily such a map, we construct an intermediate map which we define by,
@@
By exactness we can see that $\text{Ker}\partial^{*}=\text{Im}j^{*}$ and since the image of $\partial^{*}$ is empty this is the only set that determines $H_{i}(\mathcal{M},\Omega_{0}^{K})$. On the other hand, for non-contractible spaces the image of $\partial^{*}$ or $\iota^{*}$ might fail to be empty and can therefore induce a change in $H_{i}(\mathcal{M},\Omega^{K}_{0})$. This precise change can be studied through the relative homology sequence in which $H_{i}(\Omega^{K}_{0})$ or $H_{i-1}(\Omega^{K}_{0})$ is non-trivial.
@@ 3.1 Narrow and wide networks have a different impact on manifold topology
A) The average minimum rank region observed in 100 neural networks with Kaiming initialization increases with the width of the output ($n_{1}$) and is bounded by the input ($m$) layer width. Colors encode the size of the input layer with blue (m=2), light blue (m=5) orange (m=10), light orange (m=25), green (m=50), and light green(m=100). Networks with ranks below the punctured lines, matched to the colors of the input dimension, contain topologically destructive regions. B) The minimal rank decreases as a function of the mean $\mu$ of the weight distribution (normal with $\sigma=0.1$). The input dimension here is fixed at $m=10$ (black line). The colors denote networks with different output widths $n_{1}$ following the same sequence as in panel A). C) The minimal rank region of a neural network with 25 input neurons (cyan) and 50 input neurons (orange) with strictly positive weights multiplied by a rank selection matrix scales sublinearly as a function of the index of $R$. The two black curves show the bounding functions $\min\{R_{\text{indx}},n_{1}-R_{\text{indx},m}\}$ for the two corresponding networks. D) The minimal (blue) and maximal (red) ranks of a balanced connectivity (inhibitory connections are 4 times as strong as excitatory ones) network following Dale’s principle with 50 input and 100 output neurons, evaluated over strictly positive samples. The region highlighted in purple represents the regime in which there are both topologically destructive as well as topologically preserving regions.
@@ 4 How topologically destructive are neural networks after training?
For better visibility, we computed the codeword size instead of the rank. With this measure any codeword size under 20 (which is the dimension of the input) is guaranteed to be topologically destructive, whereas entries higher or equal to that are likely to not be. As can be seen in panel C of Figure 3, the lower discretizations, $f_{2}$ and $f_{3}$ specifically, which are more similar to classification tasks, on average had regions of lower codeword size compared to the pre-learning baseline. However, as one gets closer to the continuous regression version $f_{\infty}$ of this task, the codeword size of the regions was actually increased on average. This result supports the idea that neural networks solve classification tasks by orienting samples in topologically destructive regions. Regression tasks, on the other hand, are better solved in a higher rank regime, where smoothness is desired. We also explore the rank distributions of several other standard high-dimensional optimization functions taken from http://www.sfu.ca/ ssurjano [27] in the appendix.
@@
) The top row shows histograms of the codeword ranks for MNIST samples before (blue) and after (orange) training. The two vertical black lines show the intrinsic dimensionality of the MNIST dataset estimated by [21]. The row beneath shows the same, but for random noise samples instead of MNIST digits. The histogram before training is colored in green, whereas after training it is colored in purple. B) A drop in generalization, quantified in terms of the cross-entropy loss on the test data, corresponds to a reduction in the rank of the average region in layer 4. Since the true dimensionality of MNIST is not known, we highlight the region where topological destruction might start (7-13) according to the dimension estimate in [21]. As training proceeds the mean rank converges towards the lower bound of this estimate, indicating that in this case optimal performance is achieved by identifying topologically destructive regions. C) The average rank (orange) increases as a function of the number of values in the output. Plots of two dimensional functions discretized at different levels of resolution. Stars indicate a Bonferroni corrected significant difference (at $p<0.001$) between post and pretraining distributions using a Wilcoxon rank-sum test.
@@ 5 Discussion
The minimum rank region of a neural network can be manipulated by changing the sign of the rows of the weight matrix. On the other hand changing the sing of the columns, preserves Dale’s principle and in a balanced regime, guarantees that there will be both topologically destructive as well as preserving regions. Such connectivity patterns, have also been associated with optimal distance-preserving low-dimensional projections [1]. We believe that another fascinating future avenue of research would be to explore balanced rate-based recurrent neural networks obeying Dale’s principle. Such networks, could potentially provide us with insight about how neural representations change through time and thus could correspond to different stages of processing during the solution of a task.
@@ 6 Acknowledgements
This research was funded by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement Nº 945371, the Research Council of Norway Grant 300504 and the University of Oslo.

### v2

@@ Abstract
Neural networks can be thought of as applying a transformation to an input dataset. The way in which they change the topology of such a dataset often holds practical significance for many tasks, particularly those demanding non-homeomorphic mappings for optimal solutions, such as classification problems. In this work, we leverage the fact that neural networks are equivalent to continuous piecewise-affine maps, whose local rank can be used to pinpoint regions in the input space that undergo non-homeomorphic transformations, leading to alterations in the topological structure of the input dataset. Our approach enables us to make use of the relative homology sequence, with which one can study the homology groups of the quotient of a space $\mathcal{M}$ and a subset $A$, assuming some minimal properties on these spaces.
As a proof of principle, we empirically investigate the presence of low-rank (topologically-destructive) affine maps as a function of network width and mean weight. We show that in randomly initialized narrow networks, there will be regions in which the homology groups of a data manifold can change. As the width increases, the homology groups of the input manifold become more likely to be preserved. We end this part of our work by constructing highly non-random wide networks that do not have this property and relating this non-random regime to Dale’s principle, which is a defining characteristic of biological neural networks. Finally, we study simple feedforward networks trained on MNIST, as well as on toy classification and regression tasks, and show that networks manipulate the topology of data differently depending on the continuity of the task they are trained on.
@@ 1 Introduction
Following this, we introduce the idea of a representation classifying space of an input manifold, which makes use of the relative homology groups of the input manifold with respect to subspaces that fall in the low-rank regions of the polytope decomposition. We analytically compute some examples with very simple manifolds. This is followed up by empirically studying randomly initialized narrow and wide neural networks, for which we provide conditions on the width and the average weight that determine whether such networks will implement homeomorphic or non-homeomorphic maps. We also explore the impact on input topology in networks following variations of Dale’s principle. Finally, we examine which regime networks trained to solve regression and classification tasks converge to. We hope that this work provides an interesting perspective and opens up the path towards the application of more advanced topological methods in the study of neural networks.
@@ 2.1 Neural networks and decompositions
In this equation $J$ is a binary vector denoting the neurons in layer $K$ which are active and $S(x)$ is a vector function. The quantity $||J||_{0}$ will be referred to as the codeword size. These codewords form a decomposition of the input space at each layer of the network $\mathcal{M}=\bigsqcup\limits_{J}C_{J}^{K}$ for any $K$. We can also imagine a more global decomposition which is not layer dependent by defining $J$ over all the neurons in the network instead of layerwise. We call that the global polytope decomposition $C_{J}$. Another feature of this decomposition is that within each codeword set, we can express the map $\Phi^{K}$ as an affine map $Q_{J_{K}}(W_{K}\Phi^{K-1}(x)+b_{K})$, where $Q_{J_{K}}$ is a matrix which is 0 everywhere except on the diagonal $\text{diag}(Q_{J_{K}})=J_{K}$. In a way it essentially substitues the ReLU function at layer $K$ for a piecewise affine map. As one can see this definition is recursive and by going backwards through the first layer, it can be written as,
@@ 2.2 Non-homeomorphic regions of neural manifolds
Since full-rank affine maps between vector spaces are continuous (in both directions) and invertible bijections, they induce a homeomorphism between manifolds embedded in these spaces and are therefore not able to change the topology of a manifold. The only case in which the topology can change is when such an affine map projects subsets of the manifold to a subspace of a lower dimension than its embedding dimension. This is due to the fact that the projected subsets can no longer be bijectively embedded in this new lower-dimensional space. Such a degenerate map is non-injective and therefore non-homeomorphic. For a full proof see the appendix 7.2.
In order to classify the types of manifolds that one can get through such mappings, we focus on studying a standard topological invariant, namely homology groups [9]. Assuming one is working with nice (compact and Hausdorff spaces as shown in theorem 2.14 in [29]) input spaces, when the map $\Phi$ between manifolds fails to be injective, studying the homology groups $H(\Phi(\mathcal{M}))$ is equivalent to studying $H(\mathcal{M}/\sim_{\Phi})$, where $\sim_{\Phi}$ contains the equivalence classes induced by the equivalence relation $\Phi(x)=\Phi(y)$ and $\mathcal{M}/\sim_{\Phi}$ is equipped with the quotient topology.
Since the mapping $\Phi$ is a homeomorphism on all regions $\Omega_{n}$ of rank equal to $m$, we only need to consider what happens to regions of lower rank. Let $\Omega_{0}^{K}$ be a rank 0 subset from the input manifold which is projected to a single point. If this is a contractible space, projecting it to a point will not change any of the homology groups. Otherwise, quotiening this space out can have an impact. Both of these cases can be studied through the relative homology sequence [9], which is defined for a pair of spaces $(X,A)$ with $A\subset X$. If one has a map which is homeomorphic on all points of $\mathcal{M}$ except on $\Omega_{0}^{K}$ which is all projected to a point, then the relative homology sequence of the pair $(\mathcal{M},\Omega_{n}^{K})$ can be studied. Since $\Phi$ is not necessarily such a map, we construct an intermediate map which we define by,
@@
By exactness we can see that $\text{Ker}\partial^{*}=\text{Im}j^{*}$ and since the image of $\partial^{*}$ is empty this is the only set that determines $H_{i}(\mathcal{M},\Omega_{0}^{K})$. On the other hand, for non-contractible spaces the image of $\partial^{*}$ or $\iota^{*}$ might fail to be empty and can therefore induce a change in $H_{i}(\mathcal{M},\Omega^{K}_{0})$. This precise change can be studied through the relative homology sequence in which $H_{i}(\Omega^{K}_{0})$ or $H_{i-1}(\Omega^{K}_{0})$ are non-trivial.
@@ 3.1 Narrow and wide networks have a different impact on manifold topology
A) The average minimum rank region observed in 100 neural networks with Kaiming initialization increases with the width of the output ($n_{1}$) and is bounded by the input ($m$) layer width. Colors encode the size of the input layer with blue (m=2), light blue (m=5) orange (m=10), light orange (m=25), green (m=50), and light green(m=100). Networks with ranks below the punctured lines, matched to the colors of the input dimension, contain topologically destructive regions. B) The minimal rank decreases as a function of the mean $\mu$ of the weight distribution (normal with $\sigma=0.1$). The input dimension here is fixed at $m=10$ (black line). The colors denote networks with different output widths $n_{1}$ following the same sequence as in panel A). C) The minimal rank region of a neural network with 25 (cyan) or 50 (orange) input neurons with strictly positive weights multiplied by a rank selection matrix scales sublinearly as a function of the index of $R$. The two black curves show the bounding functions $\min\{R_{\text{indx}},n_{1}-R_{\text{indx},m}\}$ for the two corresponding networks, with $n_{1}=100$. D) The minimal (blue) and maximal (red) ranks of a balanced connectivity (inhibitory connections are 4 times as strong as excitatory ones) network following Dale’s principle with 50 input and 100 output neurons, evaluated over strictly positive samples. The region highlighted in purple represents the regime in which there are both topologically destructive as well as topologically preserving regions.
@@ 4 How topologically destructive are neural networks after training?
For better visibility, we computed the codeword size instead of the rank. With this measure any codeword size under 20 (which is the dimension of the input) is guaranteed to be topologically destructive, whereas entries higher or equal to that are not. As can be seen in panel C of Figure 3, the lower discretizations, $f_{2}$ and $f_{3}$ specifically, which are more similar to classification tasks, on average had regions of lower codeword size compared to the pre-learning baseline. However, when the continuous regression version $f_{\infty}$ of this task is reached the codeword size of the regions was actually increased on average. This result supports the idea that neural networks solve classification tasks by orienting samples in topologically destructive regions. Regression tasks, on the other hand, are better solved in a higher rank regime, where smoothness is desired. We also explore the rank distributions of several other standard high-dimensional optimization functions taken from http://www.sfu.ca/~ssurjano [27] in the appendix 7.5.
@@
) The top row shows histograms of the ranks across layers for MNIST samples before (blue) and after (orange) training. The two vertical black lines show the intrinsic dimensionality of the MNIST dataset estimated by [21]. The row beneath shows the same, but for normally distributed noise samples instead of MNIST digits before (green) and after (purple) training. B) A drop in generalization, quantified in terms of the cross-entropy loss on the test data, corresponds to a reduction in the rank of the average region in layer 4. Since the true dimensionality of MNIST is not known, we highlight the region where topological destruction might start (7-13) according to the dimension estimate in [21]. As training proceeds the mean rank converges towards the lower bound of this estimate, indicating that in this case optimal performance is achieved by identifying topologically destructive regions. C) The average codeword size increases as a function of the number of values in the output. Plots of two dimensional functions discretized at different levels of resolution. Stars indicate a Bonferroni corrected significant difference (at $p<0.001$) between post (orange) and pretraining (blue) distributions using a Wilcoxon rank-sum test.
@@ 5 Discussion
The minimum rank region of a neural network can be manipulated by changing the sign of the rows of the weight matrix. On the other hand changing the sign of the columns, preserves Dale’s principle and in a balanced regime, guarantees that there will be both topologically destructive as well as preserving regions. Such connectivity patterns, have also been associated with optimal distance-preserving low-dimensional projections [1]. We believe that another fascinating future avenue of research would be to explore balanced rate-based recurrent neural networks obeying Dale’s principle. Such networks, could potentially provide us with insight about how neural representations change through time and thus could correspond to different stages of processing during the solution of a task.
@@ 6 Acknowledgements
This research was funded by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement Nº 945371 and the University of Oslo.